### v2

# Existence of a Non-Uniquely Ergodic Interval Exchange Transformation with Flips Possessing Three Invariant Measures

Abstract.
We present the first explicit example of an interval exchange transformation with flips (FIET) possessing three distinct invariant ergodic measures. The proof is based on a generalization of M. Keane’s method, using the Rauzy induction adapted for FIETs, which contributes to the study of the ergodic properties of this class of dynamical systems.
1. Introduction
1.1. Overview of the main results
The main object of our research is interval exchange transformations with flips (FIETs). An FIET is a piecewise isometric map of an interval to itself with a finite number of jump discontinuities, where the orientation of at least one of the intervals of continuity is reversed; this property will be referred to as a flip (for more details on FIETs, see [Ng87]). FIETs are a generalization of classic interval exchange transformations (IETs).
IETs arise as the first return map to a transversal for an orientable measured foliation on an orientable surface or, for instance, as a first return map to a transversal for a billiard flow in a rational polygon. IETs were first mentioned by V. Arnold in 1963 in his work [Arn], where he considered IETs on three intervals.
IETs are very actively studied. The general case was first described by V. Oseledets in 1968 in [Odc]. M. Keane showed in [Kn75] that a typical IET is minimal, meaning the orbit of every point is dense. Later, H. Masur and W. Veech showed in [M82] and [V82] that almost all IETs are uniquely ergodic. Also, in a work by A. Katok [Kt], an upper bound on the number of invariant ergodic measures was obtained: an IET cannot have more than $m$ invariant ergodic measures, where $m$ is the genus of a surface with the property that the given IET is the first return map of the vertical flow to a horizontal transversal, i.e., $\lfloor\frac{n}{2}\rfloor$, where $n$ is the number of intervals. The first proof of the effectiveness of this bound was obtained by E. Sataev [St]. This result was also obtained by J. Fickenscher [Fk].
The dynamics of FIETs differs from the dynamics of classical IETs. For instance, in [Ng89], A. Nogueira proved that a typical FIET is not minimal. Due to this, the class of minimal FIETs remains largely unexplored. In particular, much less is known about the ergodic properties of minimal FIETs. Nevertheless, some progress has been made. The Hausdorff dimension of the set of minimal FIETs was estimated by A. Skripchenko and S. Troubetzkoy in [ST18]. Also, in [BL23], A. Linero Bas and G. Soler López established the existence of a minimal non-uniquely ergodic example of a FIET on 10 intervals with two distinct invariant ergodic measures. Subsequently, in [BL25], they provided an example on 6 intervals with two distinct invariant ergodic measures.
It is worth noting that one can construct examples of minimal non-uniquely ergodic FIETs with an arbitrary number of invariant measures via a straightforward procedure: by taking a known minimal IET with the desired number of measures and adding a small interval with a flip. Such a construction, however, typically leads to a reducible FIET, where the associated permutation is not a single cycle. The novelty of our work lies in constructing an example that is irreducible. In this paper, we present a proof of the following result:
Theorem 1.
There exists an irreducible minimal non-uniquely ergodic FIET with three distinct invariant ergodic measures.
The proof is based on a method proposed by M. Keane in [Kn77] for classical IETs, which involves estimating the measures of certain subintervals for each of the invariant measures and proving their difference. We generalize this approach to the case of FIETs. The key tool of analysis is a generalized Rauzy induction, adapted for FIETs, the structure of which inherits the main ideas of the original construction by G. Rauzy [Rz].
1.2. Applications of FIETs
The reason for the interest in studying FIETs is their connection to other areas of mathematics, particularly to translation surfaces and billiards in rational polygons. C. Danthony and A. Nogueira in [DN90] exploited linear involutions with flips, which are a generalization of FIETs, to show that almost all measured foliations on non-orientable surfaces have a compact leaf. In [HR18], P. Hubert and O. Paris-Romaskevich showed a connection between tiling billiards and FIETs, which allowed them to prove several theorems concerning triangle tiling billiards.
1.3. Acknowledgments
I am deeply grateful to A. Skripchenko for suggesting the problem and for her constant guidance and encouragement throughout the preparation of this paper. I would also like to thank S. Troubetzkoy for providing important comments. The research leading to these results has received funding from the Basic Research Program at the National Research University Higher School of Economics.
1.4. Some notations
$\Theta$:
The transition matrix corresponding to the Rauzy graph.
$x^{j}$:
A vector $x^{j}=(x^{j}_{1},\dots,x^{j}_{n})$ at step $j$, with components satisfying $\sum_{i}x_{i}^{j}=1$.
$|\mathbf{v}|$:
The $L_{1}$-norm of a vector $\mathbf{v}$, i.e., $|\mathbf{v}|=\sum_{i}|v_{i}|$.
$\overline{\Theta}$:
The normalized operator, $\overline{\Theta}=\frac{\Theta}{|\Theta|}$. When applied to a vector, it implies $\overline{\Theta}x=\frac{\Theta x}{|\Theta x|}$.
$\lambda_{k}(\cdot)$:
An invariant ergodic probability measure. A key notational convention is $\lambda_{k}(I^{k})=x^{k}$ and $\lambda_{k}(I^{k}_{i})=x^{k}_{i}$.
$I^{j},I^{j}_{i}$:
Subintervals. Specifically, $I^{j}_{i}$ denotes the $i$-th subinterval at the $j$-th step of the first return map.
2. Interval Exchange Transformations with Flips
In this section, we will present the formal definitions and tools that we will use to prove the results presented in the current paper.
2.1. Basic definitions
Definition 1.
An IET is a map $T:I\to I$ determined by a pair $(\lambda,\pi)$, where:
  • $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}_{+}^{n}$ is a vector of lengths. Let $L=\sum_{i=1}^{n}\lambda_{i}$.
  • $\pi=(\pi_{0},\pi_{1})\in\mathcal{S}_{n}\times\mathcal{S}_{n}$ is a pair of permutations on $\{1,\dots,n\}$.
The permutations define partitions of the domain and range of $T$. The interval $I$ is partitioned into $n$ consecutive subintervals, whose orderings are given by $\pi_{0}$ for the domain and $\pi_{1}$ for the range. For each label $k\in\{1,\dots,n\}$, there is a unique subinterval of length $\lambda_{k}$ in both the domain and the range partition.
The map $T$ sends the domain subinterval with label $k$ to the range subinterval with the same label $k$ by a translation.
$\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{1}$$\lambda_{2}$$\lambda_{3}$
Figure 1: An IET for $n=3$ with permutations $\pi_{0}=(1,2,3)$ and $\pi_{1}=(2,3,1)$.
Definition 2.
An IET $T$ is called an FIET if the partitioned interval not only has its subintervals rearranged but also has their orientation reversed. The map $T$ is defined by a triple $(\lambda,\pi,F)$:
  • $(\lambda,\pi)$ are as in the definition of an IET.
  • $F\subseteq\{1,\dots,n\}$ is a set of indices called the flip set. It is the set of intervals that reverse their orientation under the map $T$
$\lambda_{1}$$\lambda_{2}$$\lambda_{3}$$\lambda_{1}$$\lambda_{2}$$\lambda_{3}$
Figure 2: An FIET for $n=3$ with permutations $\pi_{0}=(1,2,3)$, $\pi_{1}=(2,3,1)$ and $F=\{1,3\}$.
2.2. Rauzy Induction
The main tool used in the proof of the paper’s results is a generalization of Rauzy induction.
Definition 3.
Rauzy induction is the main renormalization procedure that allows us to construct a new IET from an initial one, which has the exact same orbit structure but is defined on a smaller support interval. Rauzy induction was first introduced by Gerard Rauzy for IETs in his 1979 paper [Rz].
Right Rauzy induction: Consider an IET $T:I\to I$ on $n$ subintervals of $I=[0,1)$. We take the rightmost subintervals in the domain and range of the map. Then we induce $T$ on $0,1-\lambda)$, where $\lambda=\min(\lambda_{a_{n}},\lambda_{a_{i}})$ is the length of the shorter of the two rightmost subintervals of $I$ and $TI$. We will say that the longer interval is the "winner" ($W$) and the shorter one is the "loser" ($L$).
We adapt Rauzy induction for FIETs; it will act slightly differently on our defined pair of permutations: The induction action transforms $(\pi_{0},\pi_{1},F)$ into $(\pi^{\prime}_{0},\pi^{\prime}_{1},F^{\prime})$.
Case (a): $\boldsymbol{\lambda_{\pi_{0}(n)}>\lambda_{\pi_{1}(n)}}$:
In this case, $W=\pi_{0}(n)$, $L=\pi_{1}(n)$, and $\pi^{\prime}_{0}=\pi_{0}$. Only $\pi_{1}$ changes. We find the position of the Winner in the permutation being modified: $k=\pi_{1}^{-1}(W)=\pi_{1}^{-1}(\pi_{0}(n))$.
  1. a.1)
    $W\notin F$, insert $L$ immediately after $W$, $\pi^{\prime}_{0}=\pi_{0}$, $F^{\prime}=F\Delta\{\pi_{1}(n)\}$.
    $\pi^{\prime}_{1}(i)=\begin{cases}\pi_{1}(i)&\text{if }1\leq i\leq k\\ \pi_{1}(n)&\text{if }i=k+1\\ \pi_{1}(i-1)&\text{if }k+2\leq i\leq n\end{cases}$
  2. a.2)
    $W\in F$ , insert $L$ immediately before $W$, $\pi^{\prime}_{0}=\pi_{0}$, $F^{\prime}=F\Delta\{\pi_{1}(n)\}$.
    $\pi^{\prime}_{1}(i)=\begin{cases}\pi_{1}(i)&\text{if }1\leq i\leq k-1\\ \pi_{1}(n)&\text{if }i=k\\ \pi_{1}(i-1)&\text{if }k+1\leq i\leq n\end{cases}$
Case (b): $\boldsymbol{\lambda_{\pi_{1}(n)}>\lambda_{\pi_{0}(n)}}$:
In this case, $W=\pi_{1}(n)$, $L=\pi_{0}(n)$, and $\pi^{\prime}_{1}=\pi_{1}$. Only $\pi_{0}$ changes. We find the position of the Winner in the permutation being modified: $k=\pi_{0}^{-1}(W)=\pi_{0}^{-1}(\pi_{1}(n))$.
  1. b.1)
    $W\notin F$, insert $L$ immediately after $W$, $\pi^{\prime}_{1}=\pi_{1}$, $F^{\prime}=F\Delta\{\pi_{0}(n)\}$.
    $\pi^{\prime}_{0}(i)=\begin{cases}\pi_{0}(i)&\text{if }1\leq i\leq k\\ \pi_{0}(n)&\text{if }i=k+1\\ \pi_{0}(i-1)&\text{if }k+2\leq i\leq n\end{cases}$
  2. b.2)
    $W\in F$, insert $L$ immediately before $W$, $\pi^{\prime}_{1}=\pi_{1}$, $F^{\prime}=F\Delta\{\pi_{0}(n)\}$.
    $\pi^{\prime}_{0}(i)=\begin{cases}\pi_{0}(i)&\text{if }1\leq i\leq k-1\\ \pi_{0}(n)&\text{if }i=k\\ \pi_{0}(i-1)&\text{if }k+1\leq i\leq n\end{cases}$
These four systems completely and formally describe our new, more complex induction operation. We will denote the induction as $R_{\gamma}$, where $\gamma$ is a sequence of induction steps.
Let’s write down the 4 possible induction steps in a diagram:
b.1) If $a_{n}\notin F$, then the set $F$ does not change, and the diagram is as follows:
[…]
b.2) If $a_{n}\in F$, then the set $F$ becomes $F\Delta\{a_{i}\}$, and the diagram is as follows:
[…]
a.1) If $a_{i}\notin F$, then the set $F$ does not change, and the diagram is as follows:
$\begin{Bmatrix}a_{1}&...&a_{i}&a_{i+1}&...&a_{n}\\ ...&...&&&...&a_{i}\end{Bmatrix}$$\begin{Bmatrix}a_{1}&...&a_{i}&a_{n}&a_{i+1}&...\\ ...&...&&&...&a_{i}\end{Bmatrix}$a
a.2) If $a_{i}\in F$, then the set $F$ becomes $F\Delta\{a_{n}\}$, and the diagram is as follows:
$\begin{Bmatrix}a_{1}&...&a_{i}&a_{i+1}&...&a_{n}\\ ...&...&&&...&a_{i}\end{Bmatrix}$$\begin{Bmatrix}a_{1}&...&a_{n}&a_{i}&a_{i+1}&...\\ ...&...&&&...&a_{i}\end{Bmatrix}$a
Also, figures illustrating the generalized Rauzy induction are presented below, divided into two groups: with and without changes.
b.1a.1
Figure 3: Steps when F does not change.
b.2a.2
Figure 4: Steps when F changes.
Now we will define the transition matrix for each of these actions. Note that the old lengths are expressed in terms of the new ones according to the following principle: all interval lengths remain the same, except for the original winner’s interval, whose length becomes the sum of the new winner’s and loser’s interval lengths. The matrix for each action will look as follows:
$E_{i,j}=\begin{cases}1,&\text{if }i=j\ \text{or}\ (i=w\ \text{and}\ j=l,\ l\neq w),\\ 0,&\text{otherwise}\end{cases}$
3. Main Result
3.1. Construction of the non-uniquely ergodic example
Consider the following permutation and set $F$, which define a FIET:
$\pi=\begin{Bmatrix}\pi_{0}\\ \pi_{1}\end{Bmatrix}=\begin{Bmatrix}1&2&3&4&5&6&7&8\\ 4&5&6&7&2&1&8&3\end{Bmatrix},\ F=\{2,3,4,5,6,7\}$
Let $p_{1},p_{2},p_{3},p_{4},p_{5}>0$ and define the following Rauzy paths:
$\gamma_{1}=aaaabbabbab,$
$\gamma_{2}(p_{1})=a^{p_{1}},$
$\gamma_{3}=ba,$
$\gamma_{4}(p_{2})=b^{p_{2}},$
$\gamma_{5}=abb,$
$\gamma_{6}(p_{3})=a^{p_{3}},$
$\gamma_{7}=baaaa,$
$\gamma_{8}(p_{4})=b^{p_{4}},$
$\gamma_{9}=abbabb,$
$\gamma_{10}(p_{5})=a^{p_{5}},$
$\gamma_{11}=bba.$
Then the total path is:
$\gamma_{p}=\gamma_{1}\gamma_{2}(p_{1})\gamma_{3}\gamma_{4}(p_{2})\gamma_{5}\gamma_{6}(p_{3})\gamma_{7}\gamma_{8}(p_{4})\gamma_{9}\gamma_{10}(p_{5})\gamma_{11}$
$\gamma_{p}=aaaabbabbaba^{p_{1}}bab^{p_{2}}abba^{p_{3}}baaaab^{p_{4}}abbabba^{p_{5}}bba$
Following this path on the Rauzy diagram, we arrive at the following pair $(\pi,F)$:
$R_{\gamma_{p}}((\pi,F))=(\pi^{\prime},F^{\prime}),$
where
$\pi^{\prime}=\begin{Bmatrix}\pi_{0}\\ \pi_{1}\end{Bmatrix}=\begin{Bmatrix}1&4&2&3&5&6&7&8\\ 3&5&6&7&4&1&8&2\end{Bmatrix},\ F^{\prime}=\{2,3,4,5,6,7\}$
Note that we can consider $R^{ind_{2,3,4}}_{\gamma_{p}}$ - the induced action of the Rauzy induction, where we observe how the induction acts on intervals 2, 3, and 4.
Note that
$R^{ind_{2,3,4}}_{\gamma_{p}}(\begin{Bmatrix}2&3&4\\ 4&2&3\end{Bmatrix})=\begin{Bmatrix}4&2&3\\ 3&4&2\end{Bmatrix}$
This is a cyclic permutation with period 3:
$\begin{Bmatrix}2&3&4\\ 4&2&3\end{Bmatrix}$$\begin{Bmatrix}4&2&3\\ 3&4&2\end{Bmatrix}$$\begin{Bmatrix}3&4&2\\ 2&3&4\end{Bmatrix}$$R^{ind_{2,3,4}}_{\gamma_{p}}$$R^{ind_{2,3,4}}_{\gamma_{p}}$$R^{ind_{2,3,4}}_{\gamma_{p}}$
Thus, we find that the full cycle for our initial FIET $(\pi,F)$ is:
$\tilde{\gamma_{p}}=\gamma_{p}\gamma_{p}\gamma_{p},$
$R_{\tilde{\gamma_{p}}}((\pi,F))=(\pi,F).$
Let’s compute the transition matrix. We are interested in the transition matrix for $\gamma_{p}$, as it defines the dynamics of the exchange, and $\tilde{\gamma_{p}}$ is a power of this matrix.
Now let’s compute the transition matrices for each component of the path $\gamma_{p}$:
$\gamma_{1}=\begin{pmatrix}1&1&1&2&1&2&1&1\\ 0&1&0&0&0&0&0&0\\ 1&2&1&1&1&2&2&2\\ 1&1&1&1&0&1&1&1\\ 0&0&0&0&1&0&0&0\\ 0&1&0&1&0&1&0&1\\ 0&0&0&0&0&0&1&0\\ 1&1&0&0&0&0&1&1\end{pmatrix}$, $\gamma_{2}(p_{1})=\begin{pmatrix}1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&p_{1}&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\end{pmatrix}$.
$\gamma_{3}=\begin{pmatrix}1&1&1&1&1&0&0&1\\ 0&1&0&1&0&0&0&0\\ 0&0&1&0&1&0&0&0\\ 0&0&1&1&1&0&0&1\\ 0&0&0&0&1&0&0&0\\ 1&1&0&1&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&1\end{pmatrix}$, $\gamma_{4}(p_{2})=\begin{pmatrix}1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&p_{2}&1&0\\ 0&0&0&0&0&0&0&1\end{pmatrix}$.
$\gamma_{5}=\begin{pmatrix}1&1&0&1&0&0&0&0\\ 0&1&0&1&0&0&0&1\\ 0&0&1&0&0&0&0&0\\ 1&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&1&0\\ 0&0&0&0&0&0&1&0\\ 0&0&0&0&0&1&1&1\end{pmatrix}$, $\gamma_{6}(p_{3})=\begin{pmatrix}1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&1&0&0&0&0\\ 0&0&0&0&1&0&0&0\\ 0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1&0\\ 0&p_{3}&0&0&0&0&0&1\end{pmatrix}$.
Then the resulting matrix is:
$\Theta_{\gamma_{p}}=$ $\begin{pmatrix}9&8p_{3}+7&p_{1}+4&13&p_{1}+5&p_{2}+6&p_{2}+7&8\\ 1&p_{3}+1&0&2&0&0&0&1\\ 9&9p_{3}+8&p_{1}+3&14&p_{1}+4&2p_{2}+6&2p_{2}+8&9\\ 6&6p_{3}+5&3&9&3&p_{2}+4&p_{2}+5&6\\ 0&0&p_{1}&0&p_{1}+1&0&0&0\\ 4&4p_{3}+3&1&6&1&3&3&4\\ 0&0&0&0&0&p_{2}&p_{2}+1&0\\ 3&4p_{3}+3&1&5&1&p_{2}+2&p_{2}+3&4\end{pmatrix}$
In the next section, we will introduce a sequence of parameters $\{p^{i}\}_{i=1}^{\infty}=\{p_{1}^{i},p_{2}^{i},p_{3}^{i}\}_{i=1}^{\infty}$ and, based on constraints on this sequence, prove some estimates for the subintervals of each of the measures $\lambda_{2},\lambda_{5},\lambda_{7}$, where $\lambda_{j}$ corresponds to the vector of partition interval lengths as
$\lim_{m\to+\infty}\overline{\Theta}_{1}\overline{\Theta}_{2}...\overline{\Theta}_{m}e_{j},\text{ where }\Theta_{i}=\Theta_{\gamma_{p^{3i-2}}}\Theta_{\gamma_{p^{3i-1}}}\Theta_{\gamma_{p^{3i}}}.$
3.2. Proof of the estimates
Lemma 2.
The following estimates hold:
  1. (1)
    $\dfrac{\lambda_{7}(I^{j}_{7})}{\lambda_{7}(I^{j})}>\dfrac{1}{7}$
  2. (2)
    $2\dfrac{\lambda_{7}(I^{j}_{7})}{\lambda_{7}(I^{j})}>\dfrac{\lambda_{7}(I^{j}_{1})}{\lambda_{7}(I^{j})}$
  3. (3)
    $2\dfrac{\lambda_{7}(I^{j}_{7})}{\lambda_{7}(I^{j})}>\dfrac{\lambda_{7}(I^{j}_{4})}{\lambda_{7}(I^{j})}$
  4. (4)
    $2\dfrac{\lambda_{7}(I^{j}_{7})}{\lambda_{7}(I^{j})}>\dfrac{\lambda_{7}(I^{j}_{8})}{\lambda_{7}(I^{j})}$
  5. (5)
    $4\dfrac{\lambda_{7}(I^{j}_{7})}{\lambda_{7}(I^{j})}>\dfrac{\lambda_{7}(I^{j}_{3})}{\lambda_{7}(I^{j})}$
  6. (6)
    $\dfrac{\lambda_{7}(I^{j}_{7})}{\lambda_{7}(I^{j})}>\dfrac{\lambda_{7}(I^{j}_{5})}{\lambda_{7}(I^{j})}$
  7. (7)
    $\dfrac{\lambda_{7}(I^{j}_{5})}{\lambda_{7}(I^{j})}<\dfrac{1}{10}$
Proof.
From the construction, it is clear that $x_{6}^{j}>x_{2}^{j}$ and $x_{3}^{j}>x_{7}^{j}$. Assume that for step $j+1$ the required estimates hold. Note that by the induction hypothesis, the following is true:
$|\Theta x^{j+1}|=$ $\ 32x_{1}^{j+1}+(32p^{j}_{3}+27)x_{2}^{j+1}+(3p^{j}_{1}+12)x_{3}^{j+1}+49x_{4}^{j+1}$
$+(3p^{j}_{1}+15)x_{5}^{j+1}+(6p^{j}_{2}+21)x_{6}^{j+1}+(6p^{j}_{2}+27)x_{7}^{j+1}+32x_{8}^{j+1}>\frac{1}{2}p^{j}_{2}.$
Let’s prove that $x_{7}^{j}>x_{5}^{j}$. Consider the difference:
$|\Theta x^{j+1}|(x_{7}^{j}-x_{5}^{j})$ $=p^{j}_{2}x_{6}^{j+1}+(p^{j}_{2}+1)x_{7}^{j+1}-p^{j}_{1}x_{3}^{j+1}-(p^{j}_{1}+1)x_{5}^{j+1}$
$>p^{j}_{2}x_{6}^{j+1}+p^{j}_{2}x_{7}^{j+1}-p^{j}_{1}x_{3}^{j+1}-p^{j}_{1}x_{5}^{j+1}$
$>p^{j}_{2}x_{6}^{j+1}+(d-1)p^{j}_{1}x_{7}^{j+1}-p^{j}_{1}x_{3}^{j+1}.$
Using $4x_{7}>x_{3}$ and d1>4d-1>4, we get:
$(d-1)p^{j}_{1}x_{7}^{j+1}-p^{j}_{1}x_{3}^{j+1}>0.$
The proof of the inequality $4x_{7}^{j}>x_{3}^{j}$ follows from the fact that
$2x_{7}^{j+1}>x_{1}^{j+1},2x_{7}^{j+1}>x_{8}^{j+1}.$
Let’s prove that $2x_{7}^{j+1}>x_{8}^{j+1}$. Consider the difference:
$|\Theta x^{j+1}|(2x_{7}^{j}-x_{8}^{j})$ $=(p^{j}_{2}-2)x_{6}^{j+1}+(p^{j}_{2}-1)x_{7}^{j+1}-3x_{1}^{j+1}-(4p^{j}_{3}+3)x_{2}^{j+1}$
$\quad-x_{3}^{j+1}-5x_{4}^{j+1}-x_{5}^{j+1}-4x_{8}^{j+1}$
$>(p^{j}_{2}-2)x_{6}^{j+1}+(p^{j}_{2}-1)x_{7}^{j+1}-\frac{5}{d}-3x_{1}^{j+1}$
$\quad-x_{3}^{j+1}-5x_{4}^{j+1}-x_{5}^{j+1}-4x_{8}^{j+1}.$
For a sufficiently large $p^{j}_{2}$ ($p^{j}_{2}-1>30$), this expression is positive.
The estimates $2x_{7}^{j+1}>x_{4}^{j+1}$ (requires $p^{j}_{2}-3>58,d>56$) and $2x_{7}^{j+1}>x_{1}^{j+1}$ (requires $d>72$) are proved similarly.
Let’s prove that $x_{7}^{j+1}>\frac{1}{7}$. We need to prove that
$7(\Theta x^{j+1})_{7}-|\Theta x^{j+1}|>0.$
This inequality is equivalent to
$(p^{j}_{2}-21)x_{6}^{j+1}+(p^{j}_{2}-20)x_{7}^{j+1}-\frac{33}{d}-32x_{1}^{j+1}\\ -(3p^{j}_{1}+12)x_{3}^{j+1}-49x_{4}^{j+1}-(3p^{j}_{1}+15)x_{5}^{j+1}-32x_{8}^{j+1}>0,$
which holds for $d>33\cdot 7$.
Let’s prove that $x_{2}^{j}<\frac{1}{p^{j}_{1}}$. Since $|\Theta x^{j}|>2p^{j}_{1}$ and $x^{j}_{3}>x^{j}_{4}$, then
$x_{2}^{j}<\frac{x_{1}^{j+1}+x_{3}^{j+1}+x_{4}^{j+1}+x_{8}^{j+1}+\frac{p^{j}_{3}+1}{p^{j+1}_{1}}}{2p^{j}_{1}}<\frac{1+\frac{2}{d}}{2p^{j}_{1}}<\frac{1}{p^{j}_{1}}.$
Let’s prove that $x_{5}^{j+1}<\frac{1}{10}$. To do this, note that due to the relationships shown above, the following is true:
$|\Theta x^{j+1}|>\frac{1}{2}p^{j}_{2}.$
Now consider
$x_{5}^{j}<\frac{p^{j}_{1}x_{3}^{j+1}+(p^{j}_{1}+1)x_{5}^{j+1}}{\frac{1}{2}p^{j}_{2}}<\frac{4p^{j}_{1}x_{7}^{j+1}+(p^{j}_{1}+1)x_{7}^{j+1}}{\frac{1}{2}p^{j}_{2}}<\frac{5p^{j}_{1}+1}{\frac{d}{2}p^{j}_{1}}<\frac{11p^{j}_{1}}{dp_{1}}=\frac{11}{d}.$
Therefore, for $d>110$, we have $x_{5}^{j}<\frac{1}{10}$.
Lemma 3.
The following estimates hold:
  1. (1)
    $\frac{\lambda_{5}(I^{j}_{5})}{\lambda_{5}(I^{j})}>\frac{1}{4}$
  2. (2)
    $2\lambda_{5}(I^{j}_{3})>\lambda_{5}(I^{j}_{1})$
  3. (3)
    $\lambda_{5}(I^{j}_{3})+\lambda_{5}(I^{j}_{5})>\lambda_{5}(I^{j}_{1})$
  4. (4)
    $3\lambda_{5}(I^{j}_{6})+\lambda_{5}(I^{j}_{7})>\lambda_{5}(I^{j}_{4})$
  5. (5)
    $\lambda_{5}(I^{j}_{6})+\lambda_{5}(I^{j}_{7})\geq\lambda_{5}(I^{j}_{8})$
  6. (6)
    $\frac{\lambda_{5}(I^{j}_{2})}{\lambda_{5}(I^{j})}<\frac{1}{p^{j}_{1}}$
  7. (7)
    $\frac{\lambda_{5}(I^{j}_{6})}{\lambda_{5}(I^{j})}<\frac{7}{p^{j}_{1}}$
  8. (8)
    $\frac{\lambda_{5}(I^{j}_{7})}{\lambda_{5}(I^{j})}<\frac{1}{p^{j}_{1}}$
  9. (9)
    $\frac{\lambda_{5}(I^{j}_{8})}{\lambda_{5}(I^{j})}<\frac{22}{p^{j}_{1}}$
Proof.
Assume the following inequalities hold:
$x_{5}^{j+1}>\frac{1}{4},\quad 2x_{3}^{j+1}>x_{1}^{j+1},\quad x_{3}^{j+1}+x_{5}^{j+1}>x_{1}^{j+1},$
$3x_{6}^{j+1}+x_{7}^{j+1}>x_{4}^{j+1},\quad x_{6}^{j+1}+x_{7}^{j+1}\geq x_{8}^{j+1},$
$x_{2}^{j+1}<\frac{1}{p^{j}_{1}},\quad x_{6}^{j+1}<\frac{7}{p^{j}_{1}},\quad x_{7}^{j+1}<\frac{1}{p^{j}_{1}},\quad x_{8}^{j+1}<\frac{22}{p^{j}_{1}}.$
Note that $|\Theta x^{j+1}|>p^{j}_{1}$.
Similarly to the case of $\lambda_{7}$, it is proved that $x_{2}^{j}<\frac{1}{p^{j}_{1}}$.
Let’s prove the estimate for $x_{6}^{j}$:
$x_{6}^{j}<\frac{4x_{1}^{j+1}+\frac{5}{d}+x_{3}^{j+1}+6x_{4}^{j+1}+x_{5}^{j+1}+3x_{6}^{j+1}+3x_{7}^{j+1}+4x_{8}^{j+1}}{p^{j}_{1}}<\frac{6+\frac{5}{d}}{p^{j}_{1}}<\frac{7}{p^{j}_{1}}.$
Estimate for $x_{7}^{j}$:
$x^{j}_{7}<\frac{p^{j}_{2}x_{6}^{j+1}+(p^{j}_{2}+1)x_{7}^{j+1}}{p^{j}_{1}}<\frac{\frac{7}{d^{2}}+\frac{2}{d^{2}}}{p^{j}_{1}}<\frac{1}{p^{j}_{1}}.$
Given that $x_{8}\leq x_{7}+x_{6}$, we get:
$x^{j}_{8}<\frac{8}{p^{j}_{1}}.$
From the inequality $x_{4}<x_{7}+3x_{6}$, it follows that:
$x^{j}_{4}<\frac{22}{p^{j}_{1}}.$
Let’s prove that $x^{j}_{5}>\frac{1}{4}$. For this, it is sufficient to show that
$4p^{j}_{1}x_{3}^{j+1}+4(p^{j}_{1}+1)x_{5}^{j+1}>|\Theta x^{j+1}|,$
which is equivalent to the inequality
$(p^{j}_{1}-12)x_{3}^{j+1}+(p^{j}_{1}-11)x_{5}^{j+1}-32x_{1}^{j+1}-(32p^{j}_{3}+27)x_{2}^{j+1}\\ -49x_{4}^{j+1}-(6p^{j}_{2}+21)x_{6}^{j+1}-(6p^{j}_{2}+27)x_{7}^{j+1}-32x_{8}^{j+1}>0.$
The last inequality holds for $p^{j}_{1}>45$.
Lemma 4.
If $c>10$, then for all $i\in\{1,3,4,\dots,8\}$ the inequality holds:
$c\lambda_{2}(I^{j}_{2})>\lambda_{2}(I^{j}_{i})$
Proof.
The proof is based on comparing the coefficients of the components of the vector $x^{j+1}$ in the expressions for $x_{k}^{j}$. We compare the coefficient of $x_{2}^{j+1}$ in $x_{2}^{j}$, multiplied by $c$, with the sum of coefficients in other rows. The coefficient for $x_{2}^{j}$ is $p_{3}+1$. Below are the sums of coefficients:
$x_{1}:\ 8p_{3}+2p_{1}+2p_{2}+59,$
$x_{3}:\ 9p_{3}+2p_{1}+4p_{2}+61,$
$x_{4}:\ 6p_{3}+2p_{2}+41,$
$x_{5}:\ 2p_{1}+1,$
$x_{6}:\ 4p_{3}+25,$
$x_{7}:\ 2p_{2}+1,$
$x_{8}:\ 4p_{3}+2p_{2}+22.$
The largest coefficient is in the row for $x_{3}^{j}$. Since $p_{3}>2p_{1}+4p_{2}+61$, it is sufficient to take $c>10$ for the following inequality to be true
$cp_{3}+c>9p_{3}+2p_{1}+4p_{2}+61,$
therefore $cx_{2}^{j}>x_{i}^{j}$ for any $i\neq 2$. ∎
Lemma 5.
If $b>33$, then the following inequality holds:
$\frac{\lambda_{2}(I^{j}_{2})}{\lambda_{2}(I^{j})}>\frac{1}{b}$
Proof.
We need to prove that $b(\Theta x^{j+1})_{2}-|\Theta x^{j+1}|>0$. After substitution and simplification, we get:
$(b-32)p^{j}_{3}x_{2}^{j+1}-3p^{j}_{1}(x_{3}^{j+1}+x_{5}^{j+1})-6p^{j}_{2}(x_{6}^{j+1}+x_{7}^{j+1})-49>$
$\quad>(b-32)p^{j}_{3}x_{2}^{j+1}-c(6p^{j}_{1}+12p^{j}_{2})x_{2}^{j+1}-49=$
$\quad=\big((b-32)p^{j}_{3}-6c(p^{j}_{1}+2p_{2})\big)x_{2}^{j+1}-49>$
$\quad>(b-33)p^{j}_{3}x_{2}^{j+1}-49>(b-33)p^{j}_{3}\frac{1}{b}-49$
This is equivalent to the inequality
$(b-33)(p^{j}_{3}-49)>33\cdot 49$
Since $p^{j}_{3}>d^{3}>50^{3}$, it is sufficient for $b>33$. ∎
3.3. Non-unique ergodicity
Theorem 6.
Let there be a sequence of parameters $\{p_{1}^{i},p_{2}^{i},p_{3}^{i}\}_{i=1}^{\infty}$, where $p_{1}^{i+1}=dp_{3}^{i}=d^{2}p_{2}^{i}=d^{3}p_{1}^{i}$ and $p_{1}^{1}>d>50^{3}$. The matrix $\Theta_{i}=\Theta_{\gamma_{p^{3i-2}}}\Theta_{\gamma_{p^{3i-1}}}\Theta_{\gamma_{p^{3i}}}$ is the transition matrix for $\tilde{\gamma_{p}}$. Let a FIET be given as $(\pi,\alpha,F)$, where $\alpha\in\bigcap_{m=1}^{\infty}\overline{\Theta}_{1}...\overline{\Theta}_{m}v$, $v\in\Delta_{7}$, and $\pi$ and $F$ are from the beginning of the section. Then $\lambda_{2}$, $\lambda_{5}$, and $\lambda_{7}$ are pairwise distinct invariant ergodic measures, where $\lambda_{j}$ corresponds to the vector of partition interval lengths as
$\lim_{m\to+\infty}\overline{\Theta}_{1}\overline{\Theta}_{2}...\overline{\Theta}_{m}e_{j}.$
Proof.
Consider the partition of the initial interval $\alpha=(\alpha_{1},\dots,\alpha_{8})$.
Under the given constraints on the parameters $\{p_{1}^{i},p_{2}^{i},p_{3}^{i}\}$, the preceding lemmas imply estimates on the relative frequencies of visits to the subintervals for each of the measures.
For the measure $\lambda_{7}$: the frequency is $\lambda_{7}(\alpha_{7})>1/7$, while $\lambda_{7}(\alpha_{2})<1/p_{1}$ and $\lambda_{7}(\alpha_{5})<1/10$.
For the measure $\lambda_{5}$: the frequency is $\lambda_{5}(\alpha_{5})>1/4$, while $\lambda_{5}(\alpha_{2})<1/p_{1}$ and $\lambda_{5}(\alpha_{7})<1/p_{1}$.
For the measure $\lambda_{2}$: the frequency is $\lambda_{2}(\alpha_{2})>1/34$.
Let’s compare measures $\lambda_{5}$ and $\lambda_{7}$. Since they are probability measures, for them to be identical, it is necessary that $\lambda_{5}(A)=\lambda_{7}(A)$ for any measurable set $A$. However:
$\lambda_{5}(\alpha\setminus\alpha_{7})+\lambda_{7}(\alpha_{7})>\left(1-\frac{1}{p_{1}}\right)+\frac{1}{7}>1$
for a sufficiently large $p_{1}$. Similarly:
$\lambda_{7}(\alpha\setminus\alpha_{5})+\lambda_{5}(\alpha_{5})>\left(1-\frac{1}{10}\right)+\frac{1}{4}>1.$
Therefore, $\lambda_{5}\neq\lambda_{7}$.
Similarly, the measures $\lambda_{2}$ and $\lambda_{7}$ are distinct, since:
$\lambda_{7}(\alpha\setminus\alpha_{2})+\lambda_{2}(\alpha_{2})>\left(1-\frac{1}{p_{1}}\right)+\frac{1}{34}>1.$
The difference between measures $\lambda_{2}$ and $\lambda_{5}$ is verified in the same way. Thus, the measures $\lambda_{2}$, $\lambda_{5}$, and $\lambda_{7}$ are pairwise distinct. ∎
Proof of Theorem 1.
The proof follows from Theorem 6. ∎
We also formulate the following hypothesis about the bound on the number of non-uniquely ergodic invariant measures for an FIET:
Conjecture 7.
A FIET can have at most $\lfloor\frac{n}{2}\rfloor-1$ distinct invariant ergodic measures, where $n$ is the number of intervals.